With what acceleration will a horizontal force of 20 N move a 5 kg object along a horizontal floor if the coefficient of friction between the floor and the object is 0.18?
1
Expert's answer
2020-12-14T12:13:04-0500

Projections of net forces, acting on the object are:

"ox:" "F_x = F - F_f", where "F = 20 N" is given horizontal force and "F_f = \mu N" is static friction, "N" is normal force.

"oy:" "F_y = N - m g"

Since "a_y = 0", from the second equation, "N = m g", hence from the first equation: "m a_x = F_x = F - F_f = F - \mu N = F - \mu m g", from where "a_x = \frac{F}{m} - \mu g \approx 2.23 \frac{m}{s^2}".